The percentage composition of sodium phosphate as determined by analysis is 42.1% sodium, 18.9 % phosphorus and 39 % oxygen. Find the empirical formula of the compound[Na = 23, P = 31, O = 16].

A

Na3PO3

B

NaPO4

C

Na2PO4

D

Na3PO4

Grade:12th pass

1 Answers

Mayank Ranka
3 years ago

 

Element Percentage composition   Atomic weight    Relative No. of atoms    Simplest ratio
Sodium 42.1          23 42.1/23 = 1.83    1.83/0.61 = 3
Phosphorus 18.9     31 18.9/31 = 0.61 0.61/0.61 = 1
Oxygen 39.0     16 39.0/16 = 2.44 2.44/0.61 = 4
Thus empirical formula is Na33PO44.